According to the quantity theory of money, if an economy produces 5,000 units of output, its money supply equals $40,000 and the velocity of money equals one, then the price level will equal:

a.

$8.

b.

$0.13.

c.

$200.

d.

$1.25.
Answer #1

Answer

the equation of exchange is

MV=PY

M=money supply

V=velocity

P=price level

Y=real GDP

P=(MV)/Y=(40000*1)/5000

=$8

the price level is $8

Option a